OpenStudy (anonymous):

The radius of the circle whose equation in parametric form is P(x)= (acosx+bsinx , asinx -bcosx)

OpenStudy (perl):

so the graph is P(x) = ( acos x + bsin x , a sin x - b cos x )

OpenStudy (perl):

it probably makes more sense to use 't' as the parameter, but in any case

OpenStudy (perl):

let me graph it in parametric mode on my calculator

OpenStudy (anonymous):

actually it was theta , but i couldn't type the symbol so i used x

OpenStudy (perl):

im going to use 't' as the parameter , P(t)= (acost+bsint , asint -bcost)

OpenStudy (anonymous):

okay

OpenStudy (perl):

lets represent theta by t then :)

OpenStudy (perl):

so we have P(t) = (x(t), y(t)) where x(t) = a cos t + b sin t y(t) = a sin t - b cos t first impulse is to square x and y x^2 + y^2 = ( a cos t + b sin t ) ( a sin t - b cost ) = a^2 cos t sin t + a*b sin^t - a*b cos^2 t - b^2 cos t sin t

OpenStudy (perl):

that should stay a*b sin^2(t)

OpenStudy (anonymous):

yes then

OpenStudy (anonymous):

i got an idea....

OpenStudy (perl):

actually lets solve for x^2 separately

OpenStudy (perl):

x^2 = a^2 cos^2 (t) + b^2 * sin^2(t) + acos(t)*b*sint(t)

OpenStudy (anonymous):

 The next thing to do is pretty obvious